# Memory preservation and cooperative shielding in complex quantum networks

### v2

@@ Abstract
Complex quantum networks are powerful tools in the modeling of transport phenomena, particularly for biological systems, and enable the study of emergent phenomena in many-body quantum systems. High connectivity and long range interactions induce strong constraints on the system dynamics. Here, we study the transport properties of a quantum network described by the paradigmatic XXZ Hamiltonian, with non-trivial graph connectivity and topology, and long-range interaction. We show how long range interactions induce a memory preserving effects and strongly affect the spreading of the excitations due to cooperative shielding. We describe the memory-preserving effect in all-to-all connected regular networks with distance-independent couplings. Indeed the memory of the number of initially injected excitations is preserved over long times, being encoded in the number of frequencies present in the dynamics. Interestingly, we find that memory-preserving effects occur also in less regular graphs, such as quantum networks with either power-law node connectivity or complex, small-world type, architectures. We discuss the implications of these properties in biology-related problems, such as an application to Weber’s law in neuroscience, and their implementation in specific quantum technologies via biomimicry. We also show how the presence of long range interaction strongly affect the dynamics of the excitations in small-world networks and power law all-to-all coupled networks. Indeed due to cooperative shielding as the connectivity or the range of interaction increase the initial excitation spreads more slowly among the network and becomes strongly dependent on the initial conditions.
1 Introduction
@@
Dynamical processes in quantum networks are relevant for the engineering of quantum simulators and quantum computing platforms [1], for the investigation of biological media under driven-dissipative conditions [2, 3, 4], and to map complex, non-linear phenomena that do not require to be quantum via the novel approach called quantum-like paradigm [5, 6, 7]. Complex quantum networks are also becoming powerful tools for extracting relevant information from many-body quantum systems [8], such as emergent entanglement structures, topological instabilities, and self-similarity in quantum states, which may not be apparent by traditional methods [9, 10, 11].
@@
The importance of quantum networks, regular or complex, is enhanced by their accurate and precise engineering in current experimental setups, constituting a predominant area of development for quantum simulators [12, 13, 14], with widely transversal applications since the first proposals for quantum analog models more than 20 years ago [1]. Importantly, the current degree of experimental tunability and microscopic understanding of these platforms, both in their coherent and dissipative couplings [15], has enabled the quantum simulation of models well beyond condensed matter physics [16] or quantum computing, ranging from fundamental physics and cosmology [17], metrology and sensing [18] or – most relevant to this work – quantum thermalization and chaos [19],[20] quantum chemistry [21, 22] and biology [23].
@@
Indeed, the characterization of quantum networks dynamics under open quantum system conditions constitutes a paradigmatic testbed for understanding information processing and transport phenomena in complex media both quantum and classical [24], addressing fundamental questions in diverse contexts. These include quantum information theory [25] with transversal implications for models of quantum gravity [26, 27]; material science, in parallel with the experimental and technological development of solid-state nanodevices and cold atom technologies [28, 29]; biological media, where the role of coherent processes in out-of-equilibrium phenomena, though still debated [30], has been proven in general complex networks and biological examples [31, 32, 33, 2, 34, 24]. In fact, in more recent years the study of dynamics in complex networks topologies evolving under paradigmatic Hamiltonians, has raised attention to effectively investigate chemistry and biology applications, by different strategies that can be classified under two main categories. One is a multiscale approach that combines the quantum-chemistry computation of microscopic interaction potentials with the quantum simulation of energy or excitation transport in quantum networks [21, 23, 32]. Another one consists in mapping the complex (classical) network into a quantum network, and exploiting the underlying superposition and entanglement at work to describe the system nonlinearities [14].
@@
The quantum transport properties of a given system are governed by a set of general properties, all of which can be implemented in current quantum technology platforms: (i) the connectivity graph or network topology [35, 36], where the role of quantum interference produces drastic differences between classical and quantum regimes [24]; (ii) the presence of external couplings in the form of driving [37, 38], disorder [39, 40, 41] or dissipation [2, 42], and (iii) the coherent Hamiltonian parameters governing the static properties of the model, i.e. not just the transfer rates and interaction strengths, but also the range of such couplings [43, 44, 45] with the latter being one of our focuses. Technological development, particularly in ion traps [46] and cold gases [47, 48, 49], has allowed to explore conditions where individual quantum systems are accurately prepared to couple over large distances. The long-range nature of the interactions has been studied in detail and it is known to affect the static and dynamical properties of a quantum model [45]. As to the static properties, it has been revealed to enrich the interplay between length scales in the system [50, 51, 52], leading to non-additive energy contributions [53], novel phases of matter, and distinguished changes in spin systems [54, 55, 45] or cold-atom systems [56, 57, 58] phase diagrams, than for the case of short-range interactions. As to the dynamical behavior [53, 59], long-range interactions lead to a diversity of interesting phenomena. These concern the occurrence of non-thermal topological steady states [60, 61], excitation propagation beyond the Lieb-Robinson picture [62, 63, 64, 65], new behavior in the presence of disorder [66, 67], dissipation [68], applications to quantum reservoir computing [69] or, essential for the current work, trapped dynamics due to interference in what has been denoted as cooperative shielding [70, 71].
@@
Long-range interactions have revealed to be crucial in preserving over long time both temporal and spatial memory of excitation quenches in a regular, all-to-all connected XXZ open quantum network: this model resulted from the quantum mapping of a classical neuronal complex system and demonstrated [5] to effectively describe the sense of number [72, 73], the very general ability of humans and many animals of non-cognitively counting with constant precision of about 20% up to a few hundred items in a given space or time region [72]. Given the paradigmatic nature of the XXZ model  [74], this time-space memory preserving feature can be predicted to apply to a variety of physical systems, besides being relevant for the development of quantum technology devices.
@@
Understanding the physical mechanism governing this memory-preserving property is therefore of primary importance, especially when considering the whole variability of transport conditions represented by (i)-(iii) above. To this aim, analytical descriptions are particularly valuable in combination with information that can be extracted from quantum simulating the time and spectral signals of the network dynamical observables. In this work, we first provide complete analytical understanding of this phenomenon in regular, all-to-all connected XXZ quantum networks, in whih symmetries play a pivotal role. Then, we turn to the aforementioned points (i) and (iii) to parametrize the extent to which this phenomenon survives in more general cases: we first devote the analysis to complex quantum networks described by the XXZ Hamiltonian having couplings based on a small-world architecture; secondly, we move back to completely connected quantum networks with power-law varying interactions. In both cases, cooperative shielding and long-range interactions play a fundamental role in the memory preservation effect through non-trivial modalities. This systematic study sets the stage for a deeper exploration of information and transport processing in paradigmatic XXZ quantum (complex) networks, and the possible presence of universal behavior, revealing some degree of order in disordered structures. The article is organized as follows. In Section 2, we introduce the XXZ model, analyze its symmetry structure, and exploit the latter to identify the physical mechanism for time-space memory preservation of excitations, in the light of cooperative shielding. The analytics are performed by first studying the all-to-all connected regular XX network and then adding the $z$-axis anisotropy interaction perturbatively. In Section 3, we explore the same phenomena in highly-connected random networks and, then, in all-to-all connected regular networks with power-law decaying interactions. Finally, in Section 4, we discuss the relevance of our findings for the phenomenological description of complex systems, and the perspective research questions that our work opens up.
2 Symmetry and dynamics in all-to-all connected quantum networks via cooperative shielding
@@
The XXZ model has been studied at length for regular graphs with remarkable differences in local and global observables depending on the degree of coupling [5, 14, 70, 74]. Here, we focus on the case of highly-connected networks. We also focus on the important example of all-to-all exchange coupling (where all the $J_{ij}$ are the same independently of the distance) for different initial conditions, see Fig. 1(b)(b) Fig. 1(d)(d). The initial condition is taken to be an eigenstate of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$, which assume the general form $|\boldsymbol{\sigma}_{q}\rangle=|\sigma_{1}^{q}\dots\sigma_{L}^{q}\rangle$, with each $\sigma$ denoting either spin-up, $\uparrow$, or spin-down, $\downarrow$; this string will be dubbed as Pauli string, while the basis set $\{|\boldsymbol{\sigma}_{q}\rangle,q=1,\dots,2^{L}\}$ will be called computational basis.
In the all-to-all connected case, the interference pattern that emerges from the high connectivity leads to substantial differences in the evolution, even at long times, compared to the familiar local models. From a physical point of view, when the excitation is located at the initial site $i$, it has equal probability to tunnel to any other node if $\Delta=0$. This leads to a regular and symmetric evolution as it is seen in the power spectrum of the time signal, Fig. 1(c)(c) and $(e)$. Notably, the power spectrum shows a number of discrete frequencies equal to the number of initial excitations, see section 2.1. Notice that when considering several excitations, while the time signal profile appears complex (Fig. 1(d)(d)), the power spectrum is still well-defined (Fig. 1(e)(e)).
@@
2.1 Eigenspaces of the exchange (XX) Hamiltonian
We start by analyzing only the exchange part of the Hamiltonian (1) with all-to-all interactions, and considering it as our unperturbed mean-field component:
@@ 2.1 Eigenspaces of the exchange (XX) Hamiltonian
We show below that $\hat{\mathcal{H}}_{0}$ can be rewritten as function of the total $z$-magnetization $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$ and of the (Casimir invariant) total squared magnetization $\hat{\mathcal{M}}^{2}=\hat{\mathcal{M}}_{x}^{2}+\hat{\mathcal{M}}_{y}^{2}+\hat{\mathcal{M}}_{z}^{2}$, see [70]. We then re-cast the operators $\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}$ (equivalent for the $y$-component) in terms of $\hat{\mathcal{M}}_{z}$ and $\hat{\mathcal{M}}^{2}$ as:
$\hat{\mathcal{M}}_{x}^{2}=\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}+2\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}\,\Leftrightarrow\,\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}=\frac{\hat{\mathcal{M}}_{x}^{2}-\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}}{2}\,.$
@@
with $L$ the system size (number of spins). The result in Eq. (2) implies that although the starting Hamiltonian is the sum of two-body interacting terms, the all-to-all connectivity allows for the appearance of collective behavior, where the system behaves as if it had a single, large spin [75]. Labeling the quantum numbers associated to operators $\hat{\mathcal{M}}^{2}$ and $\hat{\mathcal{M}}_{z}$ with $l$ and $m$, respectively, we use the angular momenta composition rules for spin-$1/2$ to calculate the corresponding energy bands $E_{lm}$:
@@
$\begin{cases}f\Bigl{(}\frac{1}{2},1\Bigr{)}=1\,;\\ f(l\text{ not allowed},L)=0\,;\\ f(l,L)=f\Bigl{(}l-\frac{1}{2},L-1\Bigr{)}+f\Bigl{(}l+\frac{1}{2},L-1\Bigr{)}\,.\end{cases}$
@@
$\Delta E_{21}^{(2)}$ $=E_{2}^{(2)}-E_{1}^{(2)}=\frac{J(L-2)}{2}\,;$
$\Delta E_{32}^{(2)}$ $=E_{3}^{(2)}-E_{2}^{(2)}=\frac{JL}{2}\,;$
@@
$\langle\hat{\mathcal{O}}\rangle(t)=\langle\psi_{0}|e^{\mathrm{i}\hat{\mathcal{H}}t}\hat{\mathcal{O}}e^{-\mathrm{i}\hat{\mathcal{H}}t}|\psi_{0}\rangle=\sum_{m,n}\underbrace{\langle\psi_{0}|\phi_{m}\rangle\langle\phi_{n}|\psi_{0}\rangle\langle\phi_{m}|\hat{\mathcal{O}}|\phi_{n}\rangle}_{d_{mn}}e^{-\mathrm{i}\Delta E_{mn}t},$ (6)
where we have decomposed the initial state $|\psi_{0}\rangle$ temporal evolution in the Hamiltonian eigenbasis $\{|\phi_{n}\rangle\}$, with corresponding eigenvalues $\{E_{n}\}$ (the pedices here represent generic quantum numbers). In our specific case, the eigenstates of the system Hamiltonian $\hat{\mathcal{H}}_{0}$ are those of the angular momentum $|\alpha;l,m\rangle$, with $l$ and $m$ defined in Eq. 3 and $\alpha$ another quantum number independent of angular momentum, the initial state is a superposition of Pauli strings, and the operator $\hat{\mathcal{O}}$ is a local spin operator $\hat{\mathcal{S}}^{z}_{i}\doteq\hat{\mathcal{O}}_{0}^{(1)}$ at site $i$: this operator has the property of being the $q=0$ component of a spherical tensor of rank $r=1$, the other components being $\hat{\mathcal{O}}_{+1}^{(1)}=-c\hat{\mathcal{S}}^{+}_{i}$ and $\hat{\mathcal{O}}_{-1}^{(1)}=c\hat{\mathcal{S}}^{-}_{i}$ with $c$ a positive constant. The coefficients $\langle\psi_{0}|\phi_{m}\rangle=\langle\psi_{0}|\alpha;l,m\rangle$ and $\langle\phi_{n}|\psi_{0}\rangle=\langle\beta;l^{\prime},m^{\prime}|\psi_{0}\rangle$ are the Clebsch-Gordan coefficients between the eigenbasis and the computational basis. The matrix elements of $\hat{\mathcal{O}}_{0}^{(1)}$ can be expressed thanks to the Wigner-Eckart theorem [76] as:
@@
In other words, the system symmetries (notably invariance by rotations) guarantees that the connected matrix elements are those separated by $\Delta l\leq 1$ (and same value of $m$).
@@
This analytical result embodies the physical interpretation of the simulational findings from [5, 14]: the all-to-all connected quantum network is capable of preserving long-lasting memory in space and time of the number of initially injected excitations, this memory being encoded in frequency spectrum of the collective spin excitations.
@@
2.2 Role of the anisotropy (ZZ-term interaction)
Once the energy bands of the exchange Hamiltonian $\hat{\mathcal{H}}_{0}$ are computed, we can consider the effect of the anisotropy/interaction term in the Hamiltonian (1):

### v3

@@ Abstract
Complex quantum networks are powerful tools in the modeling of transport phenomena, particularly for biological systems, and enable the study of emergent phenomena in many-body quantum systems. High connectivity and long-range interactions induce strong constraints on the system dynamics. Here, we study the transport properties of a quantum network described by the paradigmatic XXZ Hamiltonian, with non-trivial graph connectivity and topology, and long-range interactions. We show how long-range interactions induce memory preserving effects and strongly affect the spreading of the excitations due to cooperative shielding. We describe the memory-preserving effect in all-to-all connected regular networks with distance-independent couplings. Indeed, the memory of the number of initially injected excitations is preserved over long times, encoded in the number of frequencies present in the dynamics. Interestingly, we find that memory-preserving effects occur also in less regular graphs, such as quantum networks with either power-law node connectivity or complex, small-world type, architectures. We discuss the implications of these properties in biology-related problems, such as an application to Weber’s law in neuroscience, and their implementation in specific quantum technologies via biomimicry. We also show how the presence of long-range interaction strongly affects the dynamics of the excitations in small-world networks and power law all-to-all coupled networks. Indeed, because of cooperative shielding blue, as the connectivity or the range of interaction increases, the initial excitation spreads more slowly among the network and becomes strongly dependent on the initial conditions.
preprint: arXiv:2503.05655
I Introduction
@@
Dynamical processes in quantum networks are relevant for the engineering of quantum simulators and quantum computing platforms [Daley2024], for the investigation of biological media under driven-dissipative conditions [plenio2008dephasing, Chisholm_2021, Civolani2024], and to map complex, non-linear phenomena that do not require to be quantum via the novel approach called quantum-like paradigm [YagoMalo2023, khrennikov2010paradigm, basieva2020paradigm]. Complex quantum networks are also becoming powerful tools for extracting relevant information from many-body quantum systems [LCarr], such as emergent entanglement structures, topological instabilities, and self-similarity in quantum states, which may not be apparent by traditional methods [Cirac:2010, Nokkala:2016, VParigi].
@@
The importance of quantum networks, regular or complex, is enhanced by their accurate and precise engineering in current experimental setups, constituting a predominant area of development for quantum simulators [georgescu2014simulation, Altman2024, yagomalo2024review], with widely transversal applications since the first proposals for quantum analog models more than 20 years ago [Daley2024]. Importantly, the current degree of experimental tunability and microscopic understanding of these platforms, both in their coherent and dissipative couplings [daley2014qt], has enabled the quantum simulation of models well beyond condensed matter physics [hofstetter2018condensed] or quantum computing, ranging from fundamental physics and cosmology [Bauer2023], metrology and sensing [pezze2018metro] or – most relevant to this work – quantum thermalization and chaos [eisert2010entropy],[Borgonovi2016] quantum chemistry [cao2019chemistry, Bauer2020] and biology [cao2020biology].
@@
Indeed, the characterization of quantum networks dynamics under open quantum system conditions constitutes a paradigmatic testbed for understanding information processing and transport phenomena in complex media both quantum and classical [biamonte2019complexnetworks], addressing fundamental questions in diverse contexts. These include quantum information theory [swingle2018otoc] with transversal implications for models of quantum gravity [bhattacharyya2022quantum, bentsen2019fast]; material science, in parallel with the experimental and technological development of solid-state nanodevices and cold atom technologies [datta2005transport, Lambert2021]; biological media, where the role of coherent processes in out-of-equilibrium phenomena, though still debated [Fleming2011], has been proven in general complex networks and biological examples [albert2002complexnetworks, ball2011biology, huelga2013vibrations, plenio2008dephasing, vattay2014quantum, biamonte2019complexnetworks]. In fact, in more recent years the study of dynamics in complex networks topologies evolving under paradigmatic Hamiltonians, has raised attention to effectively investigate chemistry and biology applications, by different strategies that can be classified under two main categories. One is a multiscale approach that combines the quantum-chemistry computation of microscopic interaction potentials with the quantum simulation of energy or excitation transport in quantum networks [cao2019chemistry, cao2020biology, ball2011biology]. Another one consists in mapping the complex (classical) network into a quantum network, and exploiting the underlying superposition and entanglement at work to describe the system nonlinearities [yagomalo2024review].
@@
The quantum transport properties of a given system are governed by a set of general properties, all of which can be implemented in current quantum technology platforms: (i) the connectivity graph or network topology [Kulvelis2015, mulken2016complexnetworks], where the role of quantum interference produces drastic differences between classical and quantum regimes [biamonte2019complexnetworks]; (ii) the presence of external couplings in the form of driving [Kohler2005, Alamo2024], disorder [Moix_2013, Walschaers2016, Zhou2016] or dissipation [plenio2008dephasing, Wu_2022], and (iii) the coherent Hamiltonian parameters governing the static properties of the model, i.e. not just the transfer rates and interaction strengths, but also the range of such couplings [Kloss2020, Dominguez2021, Defenu2023] with the latter being one of our focuses. Technological development, particularly in ion traps [Bruzewicz2019] and cold gases [Ritsch2013, Adams_2020, Chomaz_2023], has allowed to explore conditions where individual quantum systems are accurately prepared to couple over large distances. The long-range nature of the interactions has been studied in detail and it is known to affect the static and dynamical properties of a quantum model [Defenu2023]. As to the static properties, it has been revealed to enrich the interplay between length scales in the system [Botet:1982, Liu:2019, Defenu:2018], leading to non-additive energy contributions [Campa2009], novel phases of matter, and distinguished changes in spin systems [Maghrebi:2017, Giachetti_2021, Defenu2023] or cold-atom systems [Landig:2016, Habibian:2013, Sharma:2022] phase diagrams. As to the dynamical behavior [Campa2009, Dauxois2002], long-range interactions lead to a diversity of interesting phenomena. These concern the occurrence of non-thermal topological steady states [Borgonovi2004, Borgonovi2005], excitation propagation beyond the Lieb-Robinson picture [lieb1972bound, Schachenmayer2013, Eisert2013, Hauke2013], new behavior in the presence of disorder [Chen2023, chavez2021], dissipation [Catalano2023], applications to quantum reservoir computing [sannia2024_skineqnn] or, essential for the current work, trapped dynamics due to interference in what has been denoted as cooperative shielding [santos2016cooperative, celardo2016_shielding].
@@
The cooperative shielding effect shows how long-range interactions not always lead to a fast spreading of information, but, depending on initial conditions, the spreading of information can remain local, thus preserving the memory of the initial state for long times and affecting the transport properties of the system. Indeed, long-range interactions have revealed to be crucial in preserving over long time both temporal and spatial memory of excitation quenches in a regular, all-to-all connected XXZ open quantum network: this model resulted from the quantum mapping of a classical neuronal complex system and demonstrated [YagoMalo2023] to effectively describe the sense of number [burr2008visual, stoianov2012vision], the very general ability of humans and many animals of non-cognitively counting with constant precision of about 20% up to a few hundred items in a given space or time region [burr2008visual]. Given the paradigmatic nature of the XXZ model  [franchini2017integrable], this time-space memory preserving feature can be predicted to apply to a variety of physical systems, besides being relevant for the development of quantum technology devices.
Understanding the physical mechanism governing this memory-preserving property is therefore of primary importance, especially when considering the whole variability of transport conditions represented by (i)-(iii) above. To this aim, analytical descriptions are particularly valuable in combination with information that can be extracted from quantum simulating the time and spectral signals of the network dynamical observables.
@@
In this work we focus on two related but not identical aspects: the memory preserving effect and the transport properties of different networks in presence of long-range interactions. We first provide a complete analytical understanding of the memory effect in regular, all-to-all connected XXZ quantum networks, in which symmetries play a pivotal role. Then, we turn to the aforementioned points (i) and (iii) to parametrize the extent to which this phenomenon survives in more general cases. Specifically, we analyze complex quantum networks described by the XXZ Hamiltonian having couplings based on a small-world architecture; moreover, we move back to completely connected quantum networks with power-law varying interactions. In both cases, cooperative shielding plays a fundamental role in the memory preservation effect and in the transport properties of the system. This systematic study sets the stage for a deeper exploration of information and transport processing in paradigmatic XXZ quantum (complex) networks, and the possible presence of universal behavior, revealing some degree of order in disordered structures. The article is organized as follows. In Section II, we introduce the XXZ model, analyze its symmetry structure, and exploit the latter to identify the physical mechanism for time-space memory preservation of excitations. The analytics are performed by first studying the all-to-all connected regular XX network and then adding the $z$-axis anisotropy interaction perturbatively. In Section III, we study both transport, via the analysis of the Inverse Participation Ratio, and the memory effect in highly-connected random networks and in all-to-all connected regular networks with power-law decaying interactions. Finally, in Section IV, we discuss the relevance of our findings for the phenomenological description of complex systems, and the perspective research questions that our work opens up.
II Symmetry and dynamics in all-to-all connected quantum networks
@@
The XXZ model has been studied at length for regular graphs with remarkable differences in local and global observables depending on the degree of coupling [YagoMalo2023, yagomalo2024review, santos2016cooperative, franchini2017integrable]. Here, we focus on the case of highly-connected networks. We also focus on the important example of all-to-all exchange coupling (where all the $J_{ij}$ are the same independently of the distance) for different initial conditions, see Fig. 1(b)(b) Fig. 1(d)(d). The initial condition is taken to be an eigenstate of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$, which assume the general form $|\bm{\sigma}_{q}\rangle=|\sigma_{1}^{q}\dots\sigma_{L}^{q}\rangle$, with each $\sigma$ denoting either spin-up, $\uparrow$, or spin-down, $\downarrow$; this string will be dubbed as Pauli string, while the basis set $\{|\bm{\sigma}_{q}\rangle,q=1,\dots,2^{L}\}$ will be called computational basis.
In the all-to-all connected case, the interference pattern that emerges from the high connectivity leads to substantial differences in the evolution, even at long times, compared to the familiar local models. From a physical point of view, when the excitation is located at the initial site $i$, it has equal probability to tunnel to any other node if $\Delta=0$. This leads to a regular and symmetric evolution as it is seen in the power spectrum of the time signal, Fig. 1(c)(c) and $(e)$, generated by a standard Fast Fourier Transform (FFT) applied to the time evolution of the expectation value of local $z$-magnetization operators $\langle\hat{\mathcal{S}}_{i}^{z}\rangle$; note that this analysis is independent of the system’s dimensionality. Notably, the power spectrum shows a number of discrete frequencies equal to the number of initial excitations, see Section II.1. Notice that when considering several excitations, while the time signal profile appears complex (Fig. 1(d)(d)), the power spectrum is still well-defined (Fig. 1(e)(e)).
@@
II.1 Eigenspaces of the exchange (XX) Hamiltonian
We start by analyzing only the exchange part of the Hamiltonian of Eq.(1) with all-to-all interactions, and considering it as our unperturbed mean-field component:
@@ II.1 Eigenspaces of the exchange (XX) Hamiltonian
We show below that $\hat{\mathcal{H}}_{0}$ can be rewritten as function of the total $z$-magnetization $\hat{\mathcal{M}}_{z}=\sum_{i}\hat{\mathcal{S}}_{i}^{z}$ and of the (Casimir invariant) total squared magnetization $\hat{\mathcal{M}}^{2}=\hat{\mathcal{M}}_{x}^{2}+\hat{\mathcal{M}}_{y}^{2}+\hat{\mathcal{M}}_{z}^{2}$, see [santos2016cooperative]. We then re-cast the operators $\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}$ (equivalent for the $y$-component) in terms of $\hat{\mathcal{M}}_{z}$ and $\hat{\mathcal{M}}^{2}$ as:
$\hat{\mathcal{M}}_{x}^{2}=\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}+2\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}\,$
$\,\sum_{i<j}\hat{\mathcal{S}}_{i}^{x}\hat{\mathcal{S}}_{j}^{x}=\frac{\hat{\mathcal{M}}_{x}^{2}-\sum_{k}(\hat{\mathcal{S}}_{k}^{x})^{2}}{2}\,.$
@@
with $L$ the system size (number of spins). The result in Eq.(2) implies that although the starting Hamiltonian is the sum of two-body interacting terms, the all-to-all connectivity allows for the appearance of collective behavior, where the system behaves as if it had a single, large spin [omanakuttan2023scrambling]. Labeling the quantum numbers associated to operators $\hat{\mathcal{M}}^{2}$ and $\hat{\mathcal{M}}_{z}$ with $l$ and $m$, respectively, we use the angular momenta composition rules for spin-$1/2$ to calculate the corresponding energy bands $E_{lm}$:
@@
$\begin{cases}f\Bigl(\frac{1}{2},1\Bigr)=1\\ f(l\text{ not allowed},L)=0\\ f(l,L)=f\Bigl(l-\frac{1}{2},L-1\Bigr)+f\Bigl(l+\frac{1}{2},L-1\Bigr)\end{cases}.$
@@
$\Delta E_{21}^{(2)}$ $=E_{2}^{(2)}-E_{1}^{(2)}=\frac{J(L-2)}{2}$
$\Delta E_{32}^{(2)}$ $=E_{3}^{(2)}-E_{2}^{(2)}=\frac{JL}{2}$
@@
$\langle\hat{\mathcal{O}}\rangle(t)$ $=\langle\psi_{0}|e^{\mathrm{i}\hat{\mathcal{H}}t}\hat{\mathcal{O}}e^{-\mathrm{i}\hat{\mathcal{H}}t}|\psi_{0}\rangle$ (6)
$=\sum_{m,n}\underbrace{\langle\psi_{0}|\phi_{m}\rangle\langle\phi_{n}|\psi_{0}\rangle\langle\phi_{m}|\hat{\mathcal{O}}|\phi_{n}\rangle}_{d_{mn}}e^{-\mathrm{i}\Delta E_{mn}t},$
where we have decomposed the initial state $|\psi_{0}\rangle$ temporal evolution in the Hamiltonian eigenbasis $\{|\phi_{n}\rangle\}$, with corresponding eigenvalues $\{E_{n}\}$ (the pedices here represent generic quantum numbers). In our specific case, the eigenstates of the system Hamiltonian $\hat{\mathcal{H}}_{0}$ are those of the angular momentum $|\alpha;l,m\rangle$, with $l$ and $m$ defined in Eq.(3) and $\alpha$ another quantum number independent of angular momentum, the initial state is a superposition of Pauli strings, and the operator $\hat{\mathcal{O}}$ is a local spin operator $\hat{\mathcal{S}}^{z}_{i}\doteq\hat{\mathcal{O}}_{0}^{(1)}$ at site $i$: this operator has the property of being the $q=0$ component of a spherical tensor of rank $r=1$, the other components being $\hat{\mathcal{O}}_{+1}^{(1)}=-c\hat{\mathcal{S}}^{+}_{i}$ and $\hat{\mathcal{O}}_{-1}^{(1)}=c\hat{\mathcal{S}}^{-}_{i}$ with $c$ a positive constant. The coefficients $\langle\psi_{0}|\phi_{m}\rangle=\langle\psi_{0}|\alpha;l,m\rangle$ and $\langle\phi_{n}|\psi_{0}\rangle=\langle\beta;l^{\prime},m^{\prime}|\psi_{0}\rangle$ are the Clebsch-Gordan coefficients between the eigenbasis and the computational basis. The matrix elements of $\hat{\mathcal{O}}_{0}^{(1)}$ can be expressed thanks to the Wigner-Eckart theorem [sakurai1994] as:
$\langle\phi_{m}|\hat{\mathcal{O}}|\phi_{n}\rangle$ $=\langle\alpha;l,m|\hat{\mathcal{O}}_{0}^{(1)}|\beta;l^{\prime},m^{\prime}\rangle$
@@
In other words, the system symmetries (notably invariance by rotations) guarantee that the connected matrix elements are those separated by $\Delta l\leq 1$ (and same value of $m$).
@@
This analytical result embodies the physical interpretation of the simulational findings from [YagoMalo2023, yagomalo2024review]: the all-to-all connected quantum network is capable of preserving long-lasting memory in space and time of the number of initially injected excitations, this memory being encoded in frequency spectrum of the collective spin excitations.
@@
II.2 Role of the anisotropy (ZZ-term interaction)
Once the energy bands of the exchange Hamiltonian $\hat{\mathcal{H}}_{0}$ are computed, we can consider the effect of the anisotropy/interaction term in the Hamiltonian 1: